# Non-linear shattered pellet injection modelling in ASDEX Upgrade

### v1

@@ Abstract
Shattered pellet injection (SPI) is selected for the disruption mitigation system in ITER, due to deeper penetration, expected assimilation efficiency and prompt material delivery. This article describes non-linear simulations of SPI in the ASDEX Upgrade tokamak to test the mitigation efficiency of different injection parameters for neon-doped deuterium pellets using the JOREK code. The simulations are executed as fluid simulations. Additional marker particles are used to evolve the charge state distribution of impurities based on OpenADAS atomic data, i.e., no coronal equilibrium assumption is made. Regarding the pellet composition, neon fraction scans from 0 - 10% are performed. Numerical results show that the thermal quench (TQ) occurs in two stages. In the first stage, approximately half of the thermal energy is lost abruptly, primarily through convective and conductive transport in the stochastic fields. This stage is relatively independent of the injection parameters. In the second stage, where the majority of the remaining thermal energy is lost, radiation plays a dominant role. In cases of very low neon content, this second stage may not occur at all. A larger fraction ($\sim$20%) of the total material in the pellet is assimilated in the plasma for low neon fraction pellets (0.12%) since the full thermal collapse of the plasma occurs later than in high neon fraction scenarios. Nevertheless, the total number of assimilated neon atoms increases with increasing neon fraction. The effects of fragment size and penetration speed are further studied. Slower and smaller fragments promote edge cooling and the formation of a cold front. Faster fragments result in shorter TQ duration and higher assimilation as they reach the hotter plasma regions quicker. Using synthetic diagnostics, comparisons of general trend between simulations and experiments are conducted.
@@ 1 Introduction
Unmitigated major disruptions [1] could potentially damage the plasma facing components and the structural integrity of large fusion devices due to heat loads, electromagnetic forces, and runaway electrons (REs) [2]. To effectively mitigate all these above consequences, the shattered pellet injection (SPI) is selected as the disruption mitigation system (DMS) of ITER [3]. SPI is a technology that can actively shut down the ‘unhealthy’ plasma, by injecting a cryogenic pellet composed of pure deuterium or a mixture containing high-Z impurities [4, 5]. The injected deuterium aids in suppressing REs by increasing collisions, thereby decelerating these fast electrons, while the ablated impurities can radiate away the stored energy uniformly, hence reducing the damage to the plasma facing components. To develop a thorough understanding of SPI, essential for extrapolating its performance to future machines like ITER, SPI systems are installed on tokamak devices worldwide, including DIII-D [6], JET [7], KSTAR [8], J-TEXT [9], HL-2A [10], EAST [11], and ASDEX Upgrade (AUG) [12]. On the modelling side, many efforts are being undertaken to study the material assimilation, RE mitigation and magnetohydrodynamic (MHD) response after injection. Various codes, including the 3D MHD codes NIMROD [13], M3D-C1 [14] and JOREK [15], as well as lower dimensional codes like INDEX [16, 17] and DREAM [18], are being employed in these studies.
To validate the optimum fragment size, speed, and neon mixture fraction for the ITER DMS, the AUG tokamak has installed a flexible SPI system with three independent injecting lines, allowing for multiple simultaneous injections with options of different shatter heads [19]. After extensive lab tests and commissioning [12, 19], SPI experiments with different shattering scenarios were carried out in AUG during the 2021/2022 campaign [20]. Due to precise lab characterization of the shard cloud [21] and an excellent bolometry coverage [20], experiments in AUG and simulations for AUG scenarios can advance the understanding of the involved processes substantially and can help to robustly validate the models in view of reliable predictive simulations for ITER.
The 3D non-linear studies addressed here aim to gain a deeper understanding of complex physical phenomena by fully accounting for various self-consistent processes that are simplified or neglected in lower-dimensional models. These studies, although computationally more expensive, are crucial because they allow for the detailed inclusion of more physical processes, such as relaxation of the temperature profile by stochastic transport, current profile evolution due to resistive decay in the presence of the very large temperature variations [22], and plasmoid drift dynamics originating in the vertical motion of the electrons and ions in the inhomogeneous magnetic field [23, 24].
@@
In the present article, 3D non-linear SPI simulations for an AUG H-mode scenario are carried out using the JOREK code [29, 30]. Due to limitation of computational resource and numerical difficulty, our simulations are aimed for full TQ to early current quench (CQ) phase. Mitigation efficiency of different injection parameters for neon-doped deuterium pellets is numerically studied. Detailed analysis and variations of the injection parameters reveal fundamental dependencies. The rest of the paper will be structured as follows: In section 2, the governing reduced MHD equations are introduced, as well as the shards configuration describing the shattering process. In section 3, the effectiveness of SPI under various injection parameters are investigated. Finally, the paper is summarized and conclusions are drawn in section 4.
@@ 2.1 The governing equations
A coronal equilibrium distribution of impurity charge states had been utilized in previous JOREK SPI simulations [31]. To obtain a more accurate evolution of charge states and radiation, the impurity charge states are evolved in time self-consistently using OpenADAS atomic coefficients [32]. Here, marker particles are employed to carry the charge state evolution, while all other quantities are treated in the fluid picture. Markers are created according to the impurity ablation source $S_{imp}$ and are convected with the fluid velocity field. The ionization and recombination processes are calculated according to the local temperature and density fields. The atomic processes taking into account the charge state distribution carried by the makers provides the necessary information to calculate the radiation power density $P_{rad}$ and ionization power density $P_{ion}$.
Table 1: The SPI injection parameters for the different simulations performed in this study. Naming convention: Large/Medium/Small Fragment as well as Full/Half Velocity and the neon atomic mixture ratio in the pellet. * for the half speed case (222 m/s), to isolate the effect of the penetration speed, we maintain the same shard sizes as in LF_FV_Ne10 and MF_FV_Ne10, while manually reducing the velocities of the shards to half of their original values. It is equivalent to keeping the $v_{\perp}$ while reducing the $v_{\parallel}$ of the pellet, such that the LF_HV_Ne10 is comparable to a case with 222 m/s and 25 deg. shatter angle.
Name Ne atoms D atoms Ne frac. Speed Shatter angle Frag. Num.
LF_FV_Ne0 0 $3.0\times 10^{22}$ 0% 443 m/s 12.5 deg. 53
LF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12% 443 m/s 12.5 deg. 53
LF_FV_Ne1 $3.0\times 10^{20}$ $3.0\times 10^{22}$ 1% 443 m/s 12.5 deg. 53
LF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 443 m/s 12.5 deg. 53
MF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12% 443 m/s 20.0 deg. 199
SF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12% 443 m/s 25.0 deg. 1105
MF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 443 m/s 20.0 deg. 199
SF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 443 m/s 25.0 deg. 1105
LF_HV_Ne10* $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10% 222 m/s 25.0 deg. 53
@@
Figure 1: Shards plumes generated from the 443 m/s pellets. From left to right, the shattering angles are set as $12.5^{\circ}$ (GT3 in AUG), $20^{\circ}$ and $25^{\circ}$ (GT1 in AUG), respectively. The produced shard numbers are accordingly 53, 199 and 1105. The contours indicate the locations of different equilibrium resonant surfaces.
@@
2.2 Shards configuration
@@
The shards of the cryogenic pellet after shattering are assumed to be perfect spheres in the simulations. A rejection sampling method is applied to sequentially generate the radius of each shard, presently based on Park’s fragmentation model [33]. The sampling process is terminated once the accumulated volume of the shards is equal to that of the injected pellet. The velocities of the shards are generated based on the experimental observation that the mean shard speed tends to lie between the pellet velocity parallel to the shatter plane $v_{para}$ and the injection speed of the pellet $v_{pellet}$ [21]. For this reason, we assume the fragment velocity follows a Gaussian distribution with a mean $(v_{para}+v_{pellet})/2$ and standard deviation of 20% of the mean. The directions of the shard velocities are uniformly sampled on the unit sphere within the boundaries of a circular cone whose apex is located at the center of the sphere with a $20^{\circ}$ opening angle. The starting position of all shards is the midpoint between the bend and the exit point of the shatter tube, and the average shard direction is set parallel to the line defined by these two points and going towards the plasma.
@@
Three different shatter heads were installed on AUG during the 2021/22 experimental campaign [12, 19]. To mimic the realistic shard plume and compare the effects of different shard sizes, three different shard plumes are generated according to the above procedure, as shown in figure 1. The initial speed of the pellet is set to 443 m/s for the base scenario, with a pellet length of 10 mm and a radius of 8 mm, matching the full-sized mean pellet length around 9.6mm observed for the AUG system [19]. By varying the shattering angle from $12.5^{\circ}$ $25^{\circ}$, different shard sizes can be obtained while keeping a comparable penetration speed. The $25^{\circ}$ and $12.5^{\circ}$ cases correspond to the set-up of guide tube 1 (GT1) and GT3 in AUG, respectively. To investigate the effect of penetration speed, two additional shard configurations were created by maintaining the same shard sizes as in the 443 m/s cases with 53 and 199 shards, respectively, while manually reducing the velocities of the shards to half of their original values. All the shard configurations used in this study are listed in Tab. 1. In the simulation, each shard has its own velocity and radius, and follows an individual trajectory originating from its initial location. The velocity vector is kept constant while the radius is updated according to the ablation of the shard on its course through the plasma domain (rocket effect not taken into account).
@@
Figure 3: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity, (d) ratio of assimilated neon to total injected neon and (e) radiation power for the 53 shard cases with different neon fraction. The time of shard arrival $t_{arr}$ and exit $t_{ext}$ are marked by the purple and cyan vertical dashed lines, respectively. The time at the end of TQ is annotated by the black vertical dashed lines.
@@
Figure 4: Temporal evolution of total radiated energy, ohmic heating energy and total thermal energy for the 53 shard cases with different neon fraction. The dotted line shows the real-time radiation fraction $frad(t)=W_{rad}(t)/[W_{th}(0)-W_{th}(t)+W_{ohm}(t)]$. The percentages in the plot inform the radiation fraction before and after 1 ms (marked by vertical lines), i.e. 0 to 1 ms and 1 ms to the end of simulation. The radiation fraction before 1 ms is equal to $frad(1~{}\mathrm{ms})$, while the radiation fraction after 1 ms is calculated by $[W_{rad}(\mathrm{end})-W_{rad}(\mathrm{1~{}ms})]/[W_{th}(\mathrm{1~{}ms})-W_{th}(\mathrm{end})+W_{ohm}(\mathrm{end})-W_{ohm}(\mathrm{1~{}ms})]$.
@@
Figure 5: Poincaré plots of magnetic fields for the pure deuterium injection with 53 shards. The black markers indicate the real-time location of the shards.
@@
Figure 7: Electron temperature and density at different times (a) 0.479 ms (b) 0.497 ms (c) 0.512 ms (d) 0.540 ms during the inital thermal energy drop phase for the pure deuterium case with 53 shards. The white markers indicate the real-time location of the injected shards. The white contour illustrates the q = 3 resonant surface.
Figure 8: Flux averaged temperature profile evolution for different neon fraction during the first stage cooling.
Figure 9: Magnetic energy spectrum evolution for the 53 shard cases with different neon fraction.
Figure 10: Flux averaged temperature profile evolution for the 53 shard cases with different neon fraction. The dashed lines indicate the estimated shard trajectory.

### v2

# Non-linear MHD modelling of shattered pellet injection in ASDEX Upgrade

@@ Abstract
Shattered pellet injection (SPI) is selected for the disruption mitigation system in ITER, due to deeper penetration, expected assimilation efficiency and prompt material delivery. This article describes non-linear magnetohydrodynamic (MHD) simulations of SPI in the ASDEX Upgrade tokamak to test the mitigation efficiency of different injection parameters for neon-doped deuterium pellets using the JOREK code. The simulations are executed as fluid simulations, while additional marker particles are used to evolve the charge state distribution and radiation property of impurities based on OpenADAS atomic data, i.e., a collisional-radiative model is used. Neon fraction scans between 0 - 10% are performed. Numerical results show that the thermal quench (TQ) occurs in two stages. In the first stage, approximately half of the thermal energy is abruptly lost, primarily through convective and conductive transport in the stochastic fields. This stage is relatively independent of the neon fraction. In the second stage, where the majority of the remaining thermal energy is lost, radiation plays a dominant role. In case of pure deuterium injection, this second stage may not occur at all. A larger fraction ($\sim$20%) of the total material in the pellet is assimilated in the plasma for low neon fraction pellets ($\leq 0.12\%$) due to the full thermal collapse of the plasma occurring later than in high neon fraction scenarios. Nevertheless, the total number of assimilated neon atoms increases with increasing neon fraction. The effects of fragment size and penetration speed are then numerically studied, showing that slower and smaller fragments promote edge cooling and the formation of a cold front. Faster fragments result in shorter TQ duration and higher assimilation as they reach the hotter plasma regions quicker.
@@ 1 Introduction
Unmitigated major disruptions [1] could potentially damage the plasma facing components and the structural integrity of large fusion devices due to heat loads, electromagnetic forces, and runaway electrons (REs) [2]. To effectively mitigate all these above consequences, the ITER disruption mitigation system (DMS) is designed based on the shattered pellet injection (SPI) [3]. SPI is a technique that can actively shut down the ‘unhealthy’ plasma, by injecting a cryogenic pellet composed of pure protium or deuterium, or a mixture also containing neon [4, 5]. The injected protium/deuterium aids in suppressing REs, while the ablated impurities can radiate away the stored energy uniformly, hence reducing the localized heat loads on the plasma facing components. To develop a thorough understanding of SPI, essential for extrapolating its performance to future machines like ITER, SPI systems are installed on tokamak devices worldwide, including DIII-D [6], JET [7], KSTAR [8], J-TEXT [9], HL-2A [10], EAST [11], and ASDEX Upgrade (AUG) [12]. On the modelling side, significant efforts are being undertaken to study the material assimilation, RE mitigation and magnetohydrodynamic (MHD) response after injection. Various codes, including the 3D MHD codes NIMROD [13], M3D-C1 [14] and JOREK [15], as well as lower dimensional codes like INDEX [16, 17] and DREAM [18], are being employed in these studies.
To determine the optimum fragment size, speed, and neon mixture fraction for the ITER DMS, the AUG tokamak has installed a flexible SPI system with three independent injection lines, allowing for multiple simultaneous injections with options of different shatter heads [19]. After extensive lab tests and commissioning [12, 19], SPI experiments with different shattering scenarios were carried out in AUG during the 2022 campaign [20]. Due to precise lab characterization of the fragment cloud [21, 22] and an excellent bolometry coverage [20], experiments in AUG and simulations for AUG scenarios can substantially advance the understanding of the involved processes and can help to robustly validate the models in view of reliable predictive simulations for ITER.
The 3D non-linear MHD studies addressed here aim to gain a deeper understanding of complex physical phenomena by fully accounting for various self-consistent processes that are simplified or neglected in lower-dimensional models. These studies, although computationally more expensive, are crucial because they allow for the detailed inclusion of more physical processes, such as relaxation of the temperature profile by stochastic transport, current profile evolution due to resistive decay in the presence of the very large temperature variations [23], and plasmoid drift dynamics originating in the vertical motion of the electrons and ions in the inhomogeneous magnetic field [24, 25].
@@
In the present article, 3D non-linear SPI simulations for an AUG H-mode scenario are carried out using the JOREK code [30, 31]. Our simulations are focusing on the full thermal quench as well as the early current quench (CQ) phase. Mitigation efficiency of different injection parameters for neon-doped deuterium pellets is numerically studied. Detailed analysis and variations of the injection parameters reveal fundamental dependencies. The rest of the paper will be structured as follows: In section 2, the governing reduced MHD equations are introduced, as well as the fragment configuration describing the shattering process. In section 3, the effectiveness of SPI under various injection parameters is investigated. Comparisons of general trend between simulations and experiments are conducted using synthetic diagnostics.
@@ 2.1 The governing equations
A coronal equilibrium distribution of impurity charge states had been utilized in previous JOREK SPI simulations [32]. To obtain a more accurate evolution of charge states and radiation, the impurity charge states are self-consistently evolved in time using OpenADAS atomic coefficients [33]. Here, marker particles are employed to carry the charge state evolution, while all other quantities are treated in the fluid picture. Markers are created according to the impurity ablation source $S_{imp}$ and are convected with the fluid velocity field. The ionization and recombination processes are calculated according to the local temperature and density fields. The atomic processes taking into account the charge state distribution carried by the makers provides the necessary information to calculate the radiation power density $P_{rad}$ and ionization power density $P_{ion}$.
Table 1: The SPI injection parameters for the different simulations performed in this study. Naming convention: Large/Medium/Small Fragment as well as Full/Half Velocity and the neon atomic mixture ratio in the pellet. * for the half speed case (222 m/s), to isolate the effect of the penetration speed, we maintain the same fragment sizes as in LF_FV_Ne10 and MF_FV_Ne10, while manually reducing the velocities of the fragments to half of their original values. It is equivalent to keeping the $v_{\perp}$ while reducing the $v_{\parallel}$ of the pellet, such that the LF_HV_Ne10 is comparable to a case with 222 m/s and 25 deg. shatter angle.
Name Ne atoms D atoms Ne frac. Pre-shattered speed Shatter angle Frag. Num.
LF_FV_Ne0 0 $3.0\times 10^{22}$ 0 % 443 m/s $12.5^{\circ}$ 53
LF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12 % 443 m/s $12.5^{\circ}$ 53
LF_FV_Ne1 $3.0\times 10^{20}$ $3.0\times 10^{22}$ 1 % 443 m/s $12.5^{\circ}$ 53
LF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10 % 443 m/s $12.5^{\circ}$ 53
MF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12 % 443 m/s $20.0^{\circ}$ 199
SF_FV_Ne0.12 $3.6\times 10^{19}$ $3.0\times 10^{22}$ 0.12 % 443 m/s $25.0^{\circ}$ 1105
MF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10 % 443 m/s $20.0^{\circ}$ 199
SF_FV_Ne10 $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10 % 443 m/s $25.0^{\circ}$ 1105
LF_HV_Ne10* $3.0\times 10^{21}$ $2.7\times 10^{22}$ 10 % 222 m/s $25.0^{\circ}$ 53
@@
Figure 1: Fragments plumes generated from the 443 m/s pellets. From left to right, the shattering angles are set as $12.5^{\circ}$ (GT3 in AUG), $20^{\circ}$ and $25^{\circ}$ (GT1 in AUG), respectively. The produced fragment numbers are accordingly 53, 199 and 1105. The contours indicate the locations of different equilibrium resonant surfaces.
Figure 2: Speed and radius distributions of the fragments. The left column illustrates the distributions for the 53 fragment cases (LF_FV_Ne0, LF_FV_Ne0.12, LF_FV_Ne1 and LF_FV_Ne10), the middle column for the 199 fragment cases (MF_FV_Ne0.12 and MF_FV_Ne10), and the right column for the 1105 fragment cases (SF_FV_Ne0.12 and SF_FV_Ne10). The two half speed cases (LF_HV_Ne10 and MF_HV_Ne10) are shown in orange.
@@
Figure 4: Field aligned finite element grid including X-point used in this work. The grid resolution is 70 $\times$ 110 in the radial and poloidal direction, respectively.
2.2 Fragment configuration
@@
The fragments of the cryogenic pellet after shattering are assumed to be perfect spheres in the simulations. A rejection sampling method is applied to sequentially generate the radius of each fragment, presently based on Parks’ fragmentation model [34]. The sampling process is terminated once the accumulated volume of the fragments is equal to that of the injected pellet. The velocities of the fragments are generated based on the experimental observation that the mean fragment speed tends to lie between the pellet velocity parallel to the shatter plane $v_{para}$ and the injection speed of the pellet $v_{pellet}$ [21]. For this reason, we assume the fragment velocity follows a Gaussian distribution with a mean $(v_{para}+v_{pellet})/2$ and standard deviation of 20% of the mean. The directions of the fragment velocities are uniformly sampled on the unit sphere within the boundaries of a circular cone whose apex is located at the center of the sphere with a $20^{\circ}$ opening angle. The starting position of all fragments is the midpoint between the bend and the exit point of the shatter tube, and the average fragment direction is set parallel to the line defined by these two points and going towards the plasma.
@@
Three different shatter heads were installed on AUG during the 2022 experimental campaign [12, 19]. To mimic the realistic fragment plume and compare the effects of different fragment sizes, three different fragment plumes are generated according to the above procedure, as shown in figure 1. The pre-shattered speed of the pellet is set to 443 m/s for the base scenario, with a pellet length of 10 mm and a radius of 8 mm, matching the full-sized mean pellet length around 9.6 mm observed for the AUG system [19]. By varying the shattering angle from $12.5^{\circ}$ $25^{\circ}$, different fragment sizes can be obtained while keeping a comparable penetration speed ($cos(12.5^{\circ})=0.976$, $cos(25^{\circ})=0.906$). The $25^{\circ}$ and $12.5^{\circ}$ cases correspond to the setup of guide tube 1 (GT1) and GT3 in AUG, respectively. Although it is best to control the variable when comparing the effects of fragment size, the differences in injector geometry and trajectory were preserved in order to accurately reflect the experimental conditions for each case. To investigate the effect of penetration speed, two additional fragment configurations were created by maintaining the same fragment size distributions as in the 443 m/s cases with 53 and 199 fragments, respectively, while manually reducing the velocities of the fragments to half of their original values. This is done to separate the variables to specifically study the effect of velocity. All the fragment configurations used in this study are listed in Tab. 1, with their speed and radius distributions illustrated in figure 2. Due to the high computational cost of 3D simulations, we were unable to perform scan over different random samplings of the fragment distributions within this work. However, complementary studies using the lower dimensional model DREAM have shown that different realizations of the same statistical distribution do not significantly alter the big picture, and only have a notable impact on the disruption dynamics for trace amounts of injected neon [35]. In the simulation, each fragment has its own velocity and radius, and follows an individual trajectory originating from its initial location. The velocity vector is kept constant while the radius is updated according to the ablation of the fragment on its course through the plasma domain (rocket effect [36] not taken into account).
@@
Figure 5: Temporal evolution of (a) total thermal energy, (b) plasma current, (c) total assimilated neon quantity, (d) ratio of assimilated neon to total injected neon and (e) radiation power for the 53 fragment cases with different neon fraction. The time of fragment arrival $t_{arr}$ and exit $t_{ext}$ are marked by the purple and cyan vertical dashed lines, respectively. The times at the end of TQ are annotated by the vertical dotted lines.
@@
Figure 6: Temporal evolution of total radiated energy, ohmic heating energy and total thermal energy for the 53 fragment cases with different neon fraction. The dotted lines show the real-time radiation fraction $frad(t)=W_{rad}(t)/[W_{th}(0)-W_{th}(t)+W_{ohm}(t)]$. The vertical dashed lines indicate the time point at t = 1 ms, roughly the dividing line between the first and second stage of cooling.
@@
Figure 7: Poincaré plots of magnetic fields for the pure deuterium injection with 53 fragments. The black markers indicate the real-time location of the fragments that are not completely ablated.
@@
Figure 9: Electron temperature and density at different times (a) 0.479 ms (b) 0.497 ms (c) 0.512 ms (d) 0.540 ms during the inital thermal energy drop phase for the pure deuterium case with 53 fragments. The white markers indicate the real-time location of the injected pellet fragments. The white contour illustrates the q = 3 resonant surface.
Figure 10: Flux-averaged temperature profile evolution for different neon fractions during the first stage cooling.
Figure 11: Magnetic energy spectrum evolution for the 53 fragment cases with different neon fraction.
Figure 12: Flux-averaged temperature profile evolution for the 53 fragment cases with different neon fraction. The dashed lines indicate the estimated trajectory of the bulk fragments. For all these cases, temperature profiles are drastically degraded following the fragments arrival (0 - 0.5 ms). As the fragments penetrate inward, a localized low-temperature region is created and move inward along with the fragments (0.5 - 1.5 ms). As for the neon mixed injections, the core temperature suddenly drops to below 10 eV following the 1/1 MHD activity, and then initiating the inside-out cooling.